# Data-Locality-Aware Task Assignment and Scheduling for Distributed Job Executions

### v3

@@ Abstract
This paper investigates a data-locality-aware task assignment and scheduling problem aimed at minimizing job completion times for distributed job executions. Without prior knowledge of future job arrivals, we propose an optimal balanced task assignment algorithm (OBTA) that minimizes the completion time of each arriving job. We significantly reduce OBTA’s computational overhead by narrowing the search space of potential solutions. Additionally, we extend an approximate algorithm known as water-filling (WF) and nontrivially prove that its approximation factor equals the number of task groups in the job assignment. We also design a novel heuristic, replica-deletion (RD), which outperforms WF. To further reduce the completion time of each job, we expand the problem to include job reordering, where we adjust the order of outstanding jobs following the shortest-estimated-time-first policy. Extensive trace-driven evaluations validate the performance and efficiency of the proposed algorithms.
@@ I Introduction
Scheduling plays a critical role in the performance of big data analysis and high-performance computing, where jobs are executed across multiple distributed servers [1, 2]. These jobs typically consist of numerous tasks, each requiring access to distinct data chunks that might be replicated across different locations. Data locality, which involves assigning tasks to servers holding the required data chunks, is crucial in this context [3]. Designing online scheduling algorithms that uphold data locality in distributed job executions, without prior knowledge of job arrivals, involves two key aspects: task assignment and job reordering. Task assignment involves allocating tasks within a job to appropriate servers, conceptualized as a semi-matching problem in a bipartite graph where tasks and servers are distinct sets of nodes. Job reordering addresses the sequence in which the arrived but unfinished job tasks are executed, facilitating optimized task assignments and reducing overall job completion times.
Prior studies have addressed these issues either in isolation or in tandem [4, 5, 6, 7]. For instance, Hung et al. [4] proposed a Workload-Aware Greedy Scheduling (SWAG) strategy that schedules jobs based on their estimated completion times, prioritizing the job with the minimal completion time. Guan et al. [5] presented an offline fairness-aware task assignment algorithm that ensures max-min fairness in resource allocation across all servers. However, both [4] and [5] did not consider data replication across multiple servers, significantly reducing the complexity of the scheduling problem. Beaumont et al. [6] investigated data replication-aware scheduling, but their study focused on a single job and did not consider job reordering. Additionally, their proposed semi-matching-based strategies are only applicable to homogeneous servers with the same processing capacity and homogeneous tasks with the same execution time, which oversimplifies real-world scenarios. Guan et al. [7] studied an online scheduling problem for distributed job executions where the data chunks requested by each job could be available at multiple servers. They solved the task assignment problem using a well-constructed flow network, mapping the assignment of tasks into a maximum flow problem. For job reordering, they adopted a similar approach to [4], considering job priorities. However, they did not provide an in-depth performance analysis of the proposed algorithms.
Our work primarily focuses on the task assignment aspect of scheduling while maintaining data locality. This challenge is addressed in an online setting, where the arrival of jobs over time demands immediate scheduling decisions. The objective is to minimize the completion time of all arrived jobs, defined as the time duration from a job’s arrival to the completion of all its tasks. We study the problem in two scenarios. In the first scenario, the outstanding jobs are pending in the queues in a first-in-first-out (FIFO) manner. In the second scenario, the outstanding jobs can be prioritized and then reordered. Our contributions are summarized as follows.
@@
  • We formulate the task assignment problem as a bipartite graph-based non-linear program and develop an optimal algorithm called OBTA to solve it efficiently by significantly narrowing the search space of potential solutions.
@@
  • We extend an approximate algorithm called water-filling (WF), initially proposed for homogeneous scenarios in [7], to our settings. We nontrivially prove that WF is $K$-approximate, where $K$ is the number of task groups in a job.
  • A novel heuristic, replica-deletion (RD), is developed for task assignment. RD assigns the tasks of a job to servers with a complexity of $\mathcal{O}(M^{2}\cdot n\log n)$, where $M$ is the number of servers and $n$ is the number of tasks in a job.
@@
  • We extend WF to accommodate job reordering and introduce an early-exit technique to expedite the process, resulting in the OCWF-ACC (accelerated order-conscious WF) algorithm.
@@
  • Our trace-driven simulations underscore the effectiveness and efficiency of the proposed algorithms. OBTA, in particular, demonstrates a notable reduction in computation overhead, while WF offers a low-overhead yet effective approximation. RD further enhances performance. Additionally, compared to a job-reordered task assignment algorithm without early-exit, OCWF-ACC significantly reduces computation overhead.
The rest of this paper is organized as follows. Sec. II presents the problem formulation. Sec. III introduces the task assignment algorithms for FIFO queues. Sec. IV expands the algorithms to include job reordering. Experimental results using real job traces are discussed in Sec. V. Sec. VI reviews the related work, and Sec. VII concludes the paper.
@@ II Problem Formulation
We consider a distributed computing system consisting of $M$ servers, indexed by $m$ for $m\in\mathcal{M}:=\{1,...,M\}$. The system stores a collection of items, such as key-value pairs, objects, files, etc. They are partitioned into equally-sized data chunks. Each server holds some data chunks, and a data chunk can be replicated across multiple servers. For each $m\in\mathcal{M}$, we denote by $\mathcal{D}_{m}$ the set of data chunks stored in server $m$. The intersection $\mathcal{D}_{m}\cap\mathcal{D}_{m^{\prime}}$ can be non-empty for different $m$ and $m^{\prime}$ in $\mathcal{M}$. We make no particular assumptions about data chunk distributions. We assume that the placement of data chunks is given and does not change over time.
Jobs arrive in the system online, one by one, for execution. Each job consists of multiple independent tasks that process different data partitions. Each task takes one data chunk as input, and the tasks of the same job are assumed to have equal demands on the compute resources. We denote by $c\in\mathbb{N}^{+}$ the $c$-th job (also referred to as job $c$) in chronological order. Further, we use $\mathcal{T}_{c}$ to represent the set of tasks in job $c$, and $d_{r}$ to represent the data chunk demanded by a task $r\in\mathcal{T}_{c}$. For each task $r\in\mathcal{T}_{c}$, we define
$\mathcal{S}^{r}:=\big{\{}m\in\mathcal{M}\mid d_{r}\in\mathcal{D}_{m}\big{\}}$ (1)
as the set of its available servers. Due to data locality, each task should be assigned to one of its available servers for processing. We assume that each server $m$ has a sufficiently large queue, denoted by $q_{m}$, to store the outstanding job tasks. For analysis purposes, we divide the time into time slots of identical length.
We denote by $\mu_{m}^{c}$ the profiled number of tasks for job $c$ that can be processed at server $m$ within a time slot. We define the initial busy time of each server $m$ just before job $c$’s arrival, denoted by $b_{m}^{c}$, as the number of time slots required to finish all the outstanding job tasks in the corresponding queue. The initial busy time $b_{m}^{c}$ can be estimated using the following equation:
@@
where $o_{m}^{h}\leq|\mathcal{T}_{h}|$ is the number of outstanding tasks from job $h$ queued in $q_{m}$ just before job $c$’s arrival.
Our task is to make assignment and scheduling decisions for each arriving job. The objective is to minimize the completion time of each job. For each job, the completion time is defined as the time duration between its arrival and the completion of its last processed task. This objective is equivalent to minimizing the completion time of the last processed task issued by the job. We study the problem in two scenarios:
  • FIFO queues. The outstanding job tasks are queued in servers in a FIFO manner. The key is to assign the tasks of each arriving job to their available servers in a way that minimizes the job completion time by keeping the estimated busy times of the servers as balanced as possible.
  • Prioritized reordering. The outstanding job tasks in each queue can be reordered. The challenge lies in prioritizing the outstanding jobs in such a way that the completion times of the jobs can be further reduced.
We make no particular assumptions about job arrival times.
@@ III-A1 Programming in a Bipartite Graph
When a new job $c$ arrives, we need to assign all the tasks in $\mathcal{T}_{c}$ to appropriate available servers.
We do this by dividing these tasks into task groups, where each task group contains tasks that share the same set of available servers. We denote by $K_{c}$ the number of task groups in job $c$. Further, for each $k\in\mathcal{K}_{c}:=\{1,\ldots,K_{c}\}$, we denote by $\mathcal{T}_{c}^{k}$ and $\mathcal{S}_{c}^{k}$ the set of tasks in the $k$-th group (also referred to as group $k$) and the set of available servers for tasks in the $k$-th group, respectively. By definition, we have
@@
The task assignment problem can be modeled as a matching problem with respect to a bipartite graph. The left-side nodes in the bipartite graph are the $K_{c}$ task groups, while the right-side nodes are the $M$ servers. For each $k\in\mathcal{K}_{c}$, an edge $(k,m)$ exists if $m\in\mathcal{S}_{c}^{k}$. We can formulate the assignment problem as a non-linear integer program as follows:
@@
Recall that $b_{m}^{c}$ is the initial estimated busy time of server $m$ before the assignment of tasks in $\mathcal{T}_{c}$, calculated by (2). In $\mathcal{P}$, $n_{m}^{k}$ is the number of time slots required to process the tasks in group $\mathcal{T}_{c}^{k}$ at server $m$, given the profiled processing capacity $\mu_{m}^{c}$. Each $n_{m}^{k}$ is associated with an edge $(k,m)$ in the bipartite graph. $\Phi_{c}$ is the estimated completion time of job $c$. The first constraint in (6) means that any available server whose initial estimated busy time is larger than $\Phi_{c}$ should not participate in the assignment of tasks in $\mathcal{T}_{c}$. This constraint ensures balance among the participating servers. The second constraint in (6) ensures that all the tasks in $\mathcal{T}_{c}$ can be processed. The program $\mathcal{P}$ can be solved optimally by commercial solvers, e.g., CPLEX. [^1] https://www.ibm.com/products/ilog-cplex-optimization-studio
@@ III-A2 Narrowing the Search of $\Phi_{c}$
A commercial solver could be slow to solve the non-linear integer program in large-scale scenarios. To reduce the computational overhead, we narrow the search space of $\Phi_{c}$ by analyzing the characteristics of the problem.
After job $c$’s arrival, the number of time slots required to process all the backlogged tasks and the newly assigned tasks issued by job $c$ is upper bounded by
@@
$\Phi_{c}^{+}$ is obtained by assuming that for each available server $m$, all the tasks in $\mathcal{T}_{c}$ are assigned to it. Correspondingly, the lower bound of $\Phi_{c}$ is
@@
where $x_{k}$ is the minimum integer such that
@@
Note that $x_{k}$ is the minimal number of time slots required to process all the tasks in $\mathcal{T}_{c}^{k}$ under the assumption that group $k$ is the only task group of job $c$. With the above analysis, we can reduce the search space of $\Phi_{c}$ from $\mathbb{N}^{+}$ to the integer interval $\left[\Phi_{c}^{-},\Phi_{c}^{+}\right]$.
@@ III-A3 The Matching-based Algorithm
By introducing $\Phi_{c}^{-}$ and $\Phi_{c}^{+}$, the search space of $\Phi_{c}$ is greatly narrowed. However, $\mathcal{P}$ is still non-linear since the first constraint in (6) is piecewise. To solve $\mathcal{P}$ efficiently, we divide $\left[\Phi_{c}^{-},\Phi_{c}^{+}\right]$ into several disjoint sub-intervals based on the initial estimated busy times of available servers. The division is illustrated in Fig. 1. Specifically, we sort the initial estimated busy time $b_{m}^{c}$ of each server $m\in\bigcup_{k\in\mathcal{K}_{c}}\mathcal{S}_{c}^{k}$ in ascending order, and we denote the sorted values by $b_{1}^{\prime},b_{2}^{\prime},\ldots$. We then denote by $i$ ($j$) the minimal (maximal) index of these values such that $b_{i}^{\prime}\geq\Phi_{c}^{-}$ ($b_{j}^{\prime}\leq\Phi_{c}^{+}$). Then, we can solve $\mathcal{P}$ with the search space of $\Phi_{c}$ falling into each sub-interval $[\Phi_{c}^{-},b_{i}^{\prime}),[b_{i}^{\prime},b_{i+1}^{\prime}),\ldots,[b_{j}^{\prime},\Phi_{c}^{+}]$ in turn. Each of these problems is transformed into a linear integer program since the first constraint of (6) is no longer piecewise. The sub-intervals are checked one by one. If $\mathcal{P}$ is solvable within the current sub-interval, then its solution is the optimal one since the remaining sub-intervals cannot contain a smaller $\Phi_{c}$.
With $\mathcal{P}$ solved as detailed above, we design a task assignment algorithm named Optimal Balanced Task Assignment (OBTA). The pseudocode is presented in Algorithm 1. When the profiling of $\{\mu_{m}^{c}\}_{m,c}$ is precise, the solution of OBTA is optimal since OBTA outputs the optimal $n_{m}^{k}$’s.
@@
2       for each server $m\in\bigcup_{k\in\mathcal{K}_{c}}\mathcal{S}_{c}^{k}$ do
3             Estimate $b_{m}^{c}$ by (2)
4      Solve $\mathcal{P}$ as described in Sec. III-A3, and get the result $\{n_{m}^{k}\}_{m,k}$
5       for each task group $k\in\mathcal{K}_{c}$ do
6             for each server $m\in\mathcal{S}_{c}^{k}$ do

### v4

@@ Abstract
This paper addresses the data-locality-aware task assignment and scheduling problem for distributed job executions. Our goal is to minimize job completion times without prior knowledge of future job arrivals. We propose an Optimal Balanced Task Assignment algorithm (OBTA), which achieves minimal job completion times while significantly reducing computational overhead through efficient narrowing of the solution search space. To balance performance and efficiency, we extend the approximate Water-Filling (WF) algorithm, providing a rigorous proof that its approximation factor equals the number of task groups in a job. We also introduce a novel heuristic, Replica-Deletion (RD), which outperforms WF by leveraging global optimization techniques. To further enhance scheduling efficiency, we incorporate job ordering strategies based on a shortest-estimated-time-first policy, reducing average job completion times across workloads. Extensive trace-driven evaluations validate the effectiveness and scalability of the proposed algorithms.
@@ I Introduction
Scheduling is a fundamental component of modern service-oriented computing, enabling distributed systems to efficiently execute diverse workloads and deliver scalable services [1, 2, 3, 4]. In cloud-based environments, jobs typically comprise multiple tasks that require access to specific data chunks, which are often replicated across geographically distributed servers. Efficient data-locality-aware scheduling, which assigns tasks to servers hosting the required data chunks, is critical for reducing data transfer overhead, enhancing task execution efficiency, and ensuring compliance with stringent service-level agreements (SLAs) [5]. The design of online scheduling algorithms for such multi-task jobs introduces two central challenges: task assignment and job ordering. Task assignment entails mapping tasks to appropriate servers based on data locality, often modeled as a semi-matching problem on a bipartite graph with tasks and servers as nodes. Job ordering involves determining the execution sequence of pending tasks to optimize resource utilization and minimize job completion times, a key requirement in dynamic and resource-intensive service environments.
Existing research has made progress on these challenges, yet often under assumptions that limit applicability in real-world service systems [6, 7, 8, 9, 3, 4]. For instance, Hung et al. [6] proposed a workload-aware greedy scheduling strategy (SWAG) that prioritizes jobs based on estimated completion times. While effective in specific scenarios, this approach neglects data replication, a prevalent feature in distributed service systems. Similarly, Guan et al. [7] introduced an offline fairness-aware task assignment algorithm to ensure max-min fairness across servers, but their method lacks the adaptability needed for online and dynamic service platforms. Beaumont et al. [8] explored scheduling with data replication but constrained their analysis to single-job scenarios and homogeneous environments, which oversimplifies the heterogeneity and complexity inherent in service-oriented infrastructures. Guan et al. [9] tackled an online scheduling problem for distributed job executions, leveraging maximum flow-based techniques for task assignment. However, their work offered limited performance analysis and did not thoroughly examine job ordering strategies essential for dynamic service environments.
In this paper, we focus on developing data-locality-aware task assignment and scheduling algorithms with theoretical guarantees tailored for online distributed service environments. Our approach addresses two fundamental scenarios: (i) jobs are executed in a FIFO (first-in-first-out) manner, and (ii) jobs can be prioritized and reordered to optimize performance. By addressing these scenarios, we aim to improve the efficiency of task scheduling in service-oriented systems, tackling the dual challenges of data locality and dynamic workload management. The main contributions of this work are summarized as follows:
@@
  • We formulate the task assignment problem as a bipartite graph-based non-linear program and propose an Optimal Balanced Task Assignment algorithm (OBTA), which significantly reduces the search space and provides efficient solutions for heterogeneous service environments.
@@
  • We extend the approximate water-filling (WF) algorithm, originally proposed for homogeneous settings in [9], to heterogeneous settings. We provide a nontrivial proof that WF achieves a $K$-approximation, where $K$ represents the number of task groups in a job.
  • We propose a novel heuristic, replica-deletion (RD), for task assignment. RD generally produces better task assignments than WF.
@@
  • We extend WF to support job reordering and introduce an accelerated order-conscious WF (OCWF-ACC) algorithm, which incorporates an early-exit technique to improve computational efficiency.
@@
  • We validate the proposed algorithms through extensive trace-driven evaluations. OBTA demonstrates significant reductions in average job completion times, while WF provides a low-overhead yet effective approximation. RD enhances overall performance further. Moreover, OCWF-ACC achieves substantial reductions in computation overhead compared to job-reordered task assignment algorithms without early-exit techniques.
The remainder of this paper is organized as follows: Sec. II defines the problem formulation. Sec. III presents task assignment algorithms designed for FIFO scheduling. Sec. IV extends these algorithms to support job reordering integration. Sec. V provides experimental results based on real-world job traces, demonstrating the effectiveness of the proposed methods. Sec. VI reviews related work, highlighting the context and contributions of this study. Finally, Sec. VII concludes the paper and discusses potential future research directions.
@@ II Problem Formulation
We consider a distributed computing system comprising $M$ servers, indexed by $m$ for $m\in\mathcal{M}:=\{1,...,M\}$. The system stores a collection of data items, such as key-value pairs, objects, or files, partitioned into equally-sized data chunks. Each server holds a subset of these data chunks, and each chunk can be replicated across multiple servers. For each server $m\in\mathcal{M}$, let $\mathcal{D}_{m}$ denote the set of data chunks stored on it, where $\mathcal{D}_{m}\cap\mathcal{D}_{m’}$ can be non-empty for $m\neq m’$ ’. Key notations and acronyms used in this paper are summarized in Table I and Table II, respectively.
TABLE I: Summary of key notations.
Notation Description
$\mathcal{M}$ The set of servers
$m$ Index of servers in $\mathcal{M}$
$\mathcal{D}_{m}$ The set of data chunks in server $m\in\mathcal{M}$
$c$ Index of arriving jobs
$\mathcal{T}_{c}$ The set of tasks in job $c$
$r$ Index of tasks
$\mathcal{S}^{r}$ The available servers (due to data locality) of task $r$
$\mu_{m}^{c}$ Processing capacity of server $m$ for job $c$
$\Phi_{c}$ The estimated completion time of job $c$
$\Phi_{c}^{+}(\Phi_{c}^{-})$ The upper (lower) bound of $\Phi_{c}$ derived by OBTA
$b_{m}^{c}$ The busy time of server $m$ at the arrival of job $c$
$K_{c}$ The number of task groups in job $c$
$k$ Index of task groups
$\mathcal{T}_{c}^{k}$ The set of tasks in the $k$-th group of job $c$
$\mathcal{S}_{c}^{k}$ The set of available servers of the task group $\mathcal{T}_{c}^{k}$
ALG An algorithm ALG (and its objective value)
$\alpha_{\textsc{ALG}}$ The approximation ratio of the algorithm ALG
$I$ A job arrival instance
$\mathcal{O}_{c}$ The set of outstanding jobs at the arrival of job $c$
TABLE II: Summary of acronyms of related algorithms. OBTA-N and OBTA-P are introduced in Sec. VI.
Acronym Description
SWAG workload-aware greedy scheduling strategy
OBTA optimal balanced task assignment
NLIP non-linear integer programming
WF water-filling
RD replica deletion
OCWF order-conscious water-filling
OCWF-ACC accelerated order-conscious water-filling
LIP linearized integer programming
OBTA-N OBTA with only solution space narrowing
OBTA-P OBTA with only piecewise linearization
Jobs arrive dynamically, one at a time, for execution. Each job consists of multiple independent tasks, where each task processes a specific data partition. A task requires exactly one data chunk as input, and all tasks within a job are assumed to have identical computational demands. Let $c\in\mathbb{N}^{+}$ denote the $c$-th job (also referred to as job $c$) in chronological order. Further, let $\mathcal{T}_{c}$ represent the set of tasks in job $c$, and let $d_{r}$ denote the data chunk required by a task $r\in\mathcal{T}_{c}$. For each task $r\in\mathcal{T}_{c}$, the set of available servers capable of processing the task is defined as:
$\mathcal{S}^{r}:=\big{\{}m\in\mathcal{M}\mid d_{r}\in\mathcal{D}_{m}\big{\}}.$ (1)
To maintain data locality, each task must be assigned to one of its available servers for execution. Each server $m$ maintains a queue $q_{m}$ to store outstanding tasks, with no predefined limit on queue size. For analysis purposes, the system’s time is divided into discrete, identical time slots.
Each server $m$ has a task processing capacity characterized by $\mu_{m}^{c}$, representing the number of tasks from job $c$ that server $m$ can process within a single time slot. The initial busy time of each server $m$ just before job $c$’s arrival, denoted by $b_{m}^{c}$, is the number of time slots required to process all tasks currently queued at server $m$. This busy time can be estimated using:
@@
where $o_{m}^{h}\leq|\mathcal{T}_{h}|$ is the number of tasks from job $h$ that remain pending in $q_{m}$ when job $c$ arrives.
The objective is to design a scheduling strategy that assigns tasks to servers and determines execution priorities to minimize the completion time of each job. The completion time of a job is defined as the duration between its arrival and the completion of its last processed task, equivalent to minimizing the finish time of the last task in the job. We investigate the problem under two distinct scenarios:
  • FIFO queues. Outstanding tasks are processed in the order of their arrival (first-in-first-out). The challenge is to assign tasks of each arriving job to available servers while balancing server workloads to minimize job completion times.
  • Prioritized reordering. Outstanding tasks can be dynamically reordered in each server’s queue. This scenario focuses on prioritizing tasks to further reduce job completion times, introducing additional complexity as reordering decisions influence overall system performance.
We do not impose any specific assumptions about the timing or pattern of job arrivals.
@@ III-A1 Programming in a Bipartite Graph
When a new job $c$ arrives, the task assignment problem involves allocating all tasks in $\mathcal{T}_{c}$ to appropriate servers while ensuring data locality and minimizing the job completion time.
To facilitate this, we divide the tasks into task groups, where each task group contains tasks sharing the same set of available servers. Let $K_{c}$ denote the number of task groups for job $c$. Further, for each $k\in\mathcal{K}_{c}:=\{1,\ldots,K_{c}\}$, we denote by $\mathcal{T}_{c}^{k}$ and $\mathcal{S}_{c}^{k}$ the set of tasks in the $k$-th group (also referred to as group $k$) and the set of available servers for tasks in the $k$-th group, respectively. By definition, we have
@@
[…]
@@
[…]
@@ III-A2 Narrowing the Search of $\Phi_{c}$
Commercial solvers can be computationally expensive when solving non-linear integer programs, especially in large-scale scenarios. To mitigate this overhead, we narrow the search space of $\Phi_{c}$ by leveraging the inherent characteristics of the problem.
After the arrival of job $c$, the total number of time slots required to process both the backlogged tasks and the newly assigned tasks from job $c$ is bounded above by:
@@
This upper bound, $\Phi_{c}^{+}$, is determined under the assumption that all tasks in $\mathcal{T}_{c}$ are assigned to a single available server $m$. The corresponding lower bound of $\Phi_{c}$ is defined as:
@@
where $x_{k}$ is the minimum integer satisfying:
@@
Here, $x_{k}$ represents the minimal number of time slots required to process all tasks in $\mathcal{T}_{c}^{k}$, assuming that task group $k$ is the only group being processed for job $c$. By combining these bounds, we reduce the search space of $\Phi_{c}$ from $\mathbb{N}^{+}$ to the integer interval $\left[\Phi_{c}^{-},\Phi_{c}^{+}\right]$. This narrowing significantly reduces the computational complexity of solving the task assignment problem, making it more tractable for large-scale systems.
@@ III-A3 The Matching-based Algorithm
By introducing $\Phi_{c}^{-}$ and $\Phi_{c}^{+}$, the search space of $\Phi_{c}$ is significantly narrowed. However, the problem $\mathcal{P}$ remains non-linear due to the piecewise nature of the first constraint in (6). To address this, we divide the interval $\left[\Phi_{c}^{-},\Phi_{c}^{+}\right]$ into multiple disjoint sub-intervals based on the initial estimated busy times of the available servers. This division is illustrated in Fig. 1. Specifically, the initial estimated busy times $b_{m}^{c}$ of all servers $m\in\bigcup_{k\in\mathcal{K}_{c}}\mathcal{S}_{c}^{k}$ are sorted in ascending order, and the sorted values are denoted as $b_{1}’,b_{2}’,\ldots$ ’ , …. We then identify indices $i$ and $j$, where $i$ is the smallest index such that $b_{i}’\geq\Phi_{c}^{-}$, and $j$ is the largest index such that $b_{j}’\leq\Phi_{c}^{+}$. With this setup, the search space of $\Phi_{c}$ is divided into sub-intervals: $\left[\Phi_{c}^{-},b_{i}’\right)$ ’ ), $\left[b_{i}’,b_{i+1}’\right)$ ’ ), …, $\left[b_{j}’,\Phi_{c}^{+}\right]$. Within each sub-interval, the piecewise nature of the first constraint in (6) is resolved, transforming $\mathcal{P}$ into a linear integer program. These sub-problems are then solved sequentially. If $\mathcal{P}$ is solvable within a particular sub-interval, the corresponding solution is guaranteed to be optimal, as the remaining sub-intervals cannot yield a smaller $\Phi_{c}$.
Using this approach, we propose the Optimal Balanced Task Assignment (OBTA) algorithm. When the profiled values $\{\mu_{m}^{c}\}_{m,c}$ are accurate, OBTA produces optimal solutions by computing the exact values of $n_{m}^{k}$ for task assignment. The pseudocode for OBTA is provided in Algorithm 1.
Input: Online arriving jobs $1,2,...,c,...$, server capacities $\{\mu_{m}^{c}\}_{m,c}$
Output: Assignment solution for each task of each arriving job
@@
2 foreach $m\in\bigcup_{k\in\mathcal{K}_{c}}\mathcal{S}_{c}^{k}$ do
3      Estimate $b_{m}^{c}$ using (2)
4  Solve $\mathcal{P}$ as described in Sec. III-A3, obtaining $\{n_{m}^{k}\}_{m,k}$
5 foreach $k\in\mathcal{K}_{c}$ do
6    foreach $m\in\mathcal{S}_{c}^{k}$ do